Given the information

Product. Qty produced price. Qty produced. Price

(kg) 1995. 1995. (kg). 2000

Maize. 50 N20 100 N30

Rice. 80 N40 160 N60


Calculate:

  1. The GNP at market 2000 prices, GNP at current 1995 prices GNP at constant 1995 prices in year 2000
  2. The GNP price deflator (Nominal GNP)/Real GNP
1
Expert's answer
2021-07-12T08:44:32-0400

i.

GNP = Quantity "\times" price year

"1995 \\ (50\times 20) + (80\times 40) = N4200\\

2000 \\ (1000\times 30) + (160\times 60) =N12600\\"


ii.

GNP for 2000 year using 1995 as the base year.

GNP = Quantity "\times" price of the base year

"(1000\times 20) + (160\times 40) =N8400\\"


iii.

GNP deflator = [GNP for 2000 "\div" GNP for 2000 year using 1995 as the base year]"\times 100"

"2000\\ (\frac{12600 } {8400}) \times 100 = 150\\"